OpenStudy (anonymous):

I'll give you a medal! What is the solution for x in the proportion three over five equals twelve over x? 5 10 20 24

OpenStudy (anonymous):

Hi there! I can help you. Do you understand anything about proportions or are you completely lost?

OpenStudy (anonymous):

20 = x

OpenStudy (anonymous):

3/5 = 12/x cross multiply: 60/3 = 3x/3 reduce: 20 = x

OpenStudy (anonymous):

@ChelseaSweets We're not allowed to give just answers! You aren't helping the user understand when you do that, which is breaking OpenStudy Code of Conduct. I see that you fixed it, but next time be more careful. I don't want you to get your account suspended by a moderator! 0_0

OpenStudy (anonymous):

Ok thanks for the concern @tafkas77 ! :) (i am not saying that sarcastically i promise)

OpenStudy (anonymous):

You're welcome! I used to do it too, until I saw a lot of my friends getting suspended from their accounts... :) Thanks for not taking it the wrong way! :)

OpenStudy (anonymous):

Yeah np! I try to do the right thing i usually make them work it out then ill make them get the answer but idk why i did it differently this time.. :P

OpenStudy (anonymous):

Can you guys help me with this question right here? Estimate by rounding to the nearest whole number. twenty-three and nine-elevenths divided by three and one-seventh Answer choices: 4 6 8 twelve

OpenStudy (anonymous):

sure jus a sec

OpenStudy (anonymous):

First if i am correct u will need to turn the mixed numbers into improper fractions, do u know how to do that?

OpenStudy (anonymous):

@ChelseaSweets Sometimes I find myself just feeling lazy... :P Gotta keep working on it. @twilaswift I think Chelsea's got this one, and I have to go. But I'm sure I'm leaving you in good hands. P.S. *fan her* :D

OpenStudy (anonymous):

lol :) thanks @tafkas77

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@twilaswift do u know how to turn a mixed number into an improper fraction?